OpenStudy (anonymous):

solve for p. 3(p+q)=p my answer choices all become negative and I dont understand this at all someone please help

jimthompson5910 (jim_thompson5910):

\[\large 3(p+q)=p\] \[\large 3p+3q=p\] \[\large 3q=p-3p\] \[\large 3q=-2p\] \[\large -2p = 3q\] \[\large p = -\frac{3q}{2}\]

OpenStudy (anonymous):

wow thank you so much I get it thanks for explaining it to me!!!

jimthompson5910 (jim_thompson5910):

you're welcome